OpenStudy (anonymous):

Line A passes through the points (-2, 0) and (1, 12). Line B passes through the points (7, -2) and (-5, 22). Find the point where line A intersects line B.

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi?x%20 \%20=%20\%204,%20\%20y%20\%20=%20\%20-2

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi?x%20 \%20=%20\%20-2,%20\%20y%20\%20=%20\%20-2

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi?x%20 \%20=%20\%20\frac{32}{3},%20\%20y%20\%20=%20\%20\frac{2}{3}

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi?x%20 \%20=%20\%20\frac{2}{3},%20\%20y%20\%20=%20\%20\frac{32}{3}

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

Those answers are tiny tiny pictures, but i will tell you the answer

OpenStudy (anonymous):

What is it

OpenStudy (anonymous):

Here ill write them

OpenStudy (anonymous):

x=4,y=-2 x=-2,y=-2 x=32/3,y=2/3 x=2/3,y=32/3

OpenStudy (danjs):

k 1 sec

OpenStudy (anonymous):

k

OpenStudy (danjs):

D

OpenStudy (anonymous):

Line A has a slope of -2/3 and passes through the point https://media.studyisland.com/cgi-bin/mimetex.cgi? \left(3,%20\%209\,%20\frac{4}{5}\right) Line B has a slope of 1/3 and passes through the point (-25,-1) Find the point where line A intersects line B.

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

x=-1,y=7 x=5,y=9 x=27 1/2,y=0 x=9,y=5

OpenStudy (danjs):

can you retype that w/o the link, it didnt work

OpenStudy (danjs):

A slope -2/3 and point?

OpenStudy (anonymous):

its 3, 9 4/5

OpenStudy (anonymous):

No @DanJS the slope passes through the point

OpenStudy (danjs):

line A --- Slope (-2/3) and point (x,y) Line B --- slope (1/3) and point (-25,-1) what is that (x,y)?

OpenStudy (anonymous):

NVM i restarted it to get new ones

OpenStudy (anonymous):

Use elimination to find the solution to the system of equations.

OpenStudy (danjs):

those links don't work